Find the pH for a solution that is 0.5 M in NH4Cl and 2.0 M in NH3 when the Kb= 1.75*10-4 for NH3 . NH4Cl is a salt that ionizes to NH4+ and Cl-.
Answer #1

The accepted Kb value for NH3 is 1.75 x 10-5 but you have given the value 1.75 x 10-4 . So I will calculate the pH with your value

NH4Cl is an ionic salt which completely dissociates to NH4+ and Cl-

Concentration of NH4+ = concentration of NH4Cl

Concentration of NH4+ = 0.5 M

Kb = 1.75 x 10-4

pKb = -log(Kb)

pKb = -log(1.75 x 10-4)

pKb = 3.76

According to Henderson-Hasselbalch equation

pOH = pKb + log([conjugate acid] / [weak base])

where [conjugate acid] = concentration of conjugate acid = [NH4+] = 0.5 M

[weak base] = concentration of weak base = [NH3] = 2.0 M

pKb = 3.76

Substituting these values,

pOH = 3.76 + log(0.5 M / 2.0 M)

pOH = 3.76 - 0.60

pOH = 3.16

pH = 14 - pOH

pH = 14 - 3.16

pH = 10.84